2r+-7=1r--5
I am not sure if the negative 5 will become a positive because of the subtration in front of it. this how I would do the question
2r-1r+-7+7=1r-1r--5+7
r=12


Answer by stanbon(75887)

2r+-7=1r--5
2r-7=r+5
Subtract r from both sides to get:
r-7=5
Add 7 to both sides to get:
r=12
Cheers,
Stan H.